OpenStudy (anonymous):

The perimeter of a rectangle is 200 feet. Describe the possible lengths of a side if the area of the rectangle is not to exceed 900 square feet

OpenStudy (anonymous):

ANSWER: The length (in feet) of a side is in the interval _____ or in the interval _____

OpenStudy (dan815):

let x be the width and y be the length of a rectangle in feet the perimeter is made up of 2 lines of width and 2 of length so 2x+2y=200 now the area of the rectangle is given by x*y<900

OpenStudy (anonymous):

okay I get it so far

OpenStudy (dan815):

thinkinngg

OpenStudy (anonymous):

the answers I got were 10 and 20 but those were wrong

OpenStudy (dan815):

okay so lets say 2y=200-2x y=100-x A=x*y A=x*(100-x) A=100x-x^2 900=100x-x^2

OpenStudy (dan815):

Now you are okay with all the Xs as long as they give a number below 900

OpenStudy (dan815):

so we have 900>=100x-x^2 x^2-100x+900>=0 (x -90)(x -10)>=0 x=90 or 10 which means y=10 or 90

OpenStudy (dan815):

now we have to see if x has to be greater than 90 or 10, or if it has to be less

OpenStudy (dan815):

200>=x>=90 these are the 2 intervals that will work out 0=<x=<10

OpenStudy (dan815):

would you like to see a graphical picture of this

OpenStudy (anonymous):

How do you write these in interval notation?

OpenStudy (dan815):

umm

OpenStudy (dan815):

with brackets and stuff

OpenStudy (anonymous):

yeah

OpenStudy (dan815):

look at that btw, u can see how chosing x below 10, will result in less area, as well as chosing x greater than 90 is resulting is less area than 900

OpenStudy (dan815):